OpenStudy (anonymous):

Starting with F(x)=∫(-2 to x) 3t^2(cos(t^3)+2)dt, use the substitution u(t)=t^3 to rewrite the definite integral.

ganeshie8 (ganeshie8):

\(\large F(x) = \int \limits_{-2}^x 3t^2 (\cos(t^3) + 2)dt\)

ganeshie8 (ganeshie8):

substitute \(\large u = t^3\)

ganeshie8 (ganeshie8):

\(\large du = 3t^2 dt\)

ganeshie8 (ganeshie8):

next work the bounds, do u knw how to figure out the new bounds for \(u\) ?

OpenStudy (anonymous):

thank you for replying! I'm not completely sure, can you please help?

ganeshie8 (ganeshie8):

the original bounds are (-2, x) right ?

ganeshie8 (ganeshie8):

when u substitute u = t^3, the bounds change to : ((-2)^3, x^3) (-8, x^3)

ganeshie8 (ganeshie8):

so the integral becomes :- \(\large F(x) = \int \limits_{-8}^{x^3} (\cos(u) + 2) du\)

ganeshie8 (ganeshie8):

see if that makes some sense.. .

OpenStudy (anonymous):

aha! I didn't realize how simple that was. thank you!

ganeshie8 (ganeshie8):

:) u wlc !